How many litres of 14% alcohol solution must be mixed with 20L of a 50% solution to get a 30% solution.
How do you set this problem up. I thought .14x+.5(x+20)=.3 but I am way off.

Answer by stanbon(29603)

How many litres of 14% alcohol solution must be mixed with 20L of a 50% solution to get a 30% solution.
Let "x" be amount of 14% solution.
Then 0.14x liters of this is alcohol.
Also, 0.50(20L)= 10 liters is alcohol.
EQUATION:
alcohol + alcohol = 0.30(x+20)liters
0.14x + 10 = 0.30x + 6
Multiply through by 100 to get:
14x + 1000 = 30x + 600
-16x = -400
x= 25 liters (this is the amount of 14% solution you need to add)
Cheers,
Stan H.